### v3

@@ 1 Introduction
We present a simple LLM culture perception extraction framework that does not involve human labeling or external knowledge bases, and can be applied on any LLM and study any culture. We first use natural language prompts to elicit LLM generations on 8 culture-related topics for 110 countries and regions, using gpt-4, llama2-13b and mistral-7b. From these generation we extract culture symbols, i.e. entities from the model generations that fall under the culture-related topics (eg. “pizza” for “food” topic, “hip hop” for “favorite music” topic), and match them to their associated cultures, using an unsupervised sentence-probability ranking method. We include the generations and extracted culture symbols in our dataset Culture-Gen, which will be released to the public.
@@
In our paper, we also demonstrate how the community can use Culture-Gen for uncovering LLM’s global culture perceptions. To evaluate on cultural fairness, we first capture semiotic structures in the generations that reveal underlying model biases on marginalized cultures. We find that models tend to precede generations with the word “traditional” for cultures in Asian, Eastern European, and African-Islamic countries, as many as 30% for llama2-13b generations and almost 100% for gpt-4 generations. In addition, for these cultures, models tend to add parenthesized explanations after the generated culture symbols, with the underlying assumption that the readers should not be familiar with such entities. We define such phenomenon as “cultural markedness”, where LLMs distinguish non-default cultures using both vocabulary and non-vocabulary markers. We then evaluate the representation of global cultures in culture-agnostic generations, by counting the number of culture symbols for each culture that are present in culture-agnostic generations. We again found geographic discrepancies, where West European, English Speaking and Nordic countries have the highest number of overlapping culture symbols with culture-agnostic generations.
@@ 2 Related Work
Another line of work explores cultural diversity under multilingual settings (Mukherjee et al., 2023), incorporating geo-diverse multilingual probing Yin et al. (2022), and investigating multicultural biases using the Word Embedding Association Test (WEAT) across 24 languages (Mukherjee et al., 2023). In particular, low-resource languages have been studied as media of under-represented cultures like African American (Deas et al., 2023) and Indonesian Wibowo et al. (2023) languages and cultures. AlKhamissi et al. (2024). Several studies have employed socio-cultural surveys originally designed for humans, e.g., the World Values Survey (WVS) and Pew Global Attitudes Survey, to evaluate the cultural understanding of LLMs (Ramezani & Xu, 2023; Tao et al., 2023; Durmus et al., 2023) and extrapolate training data for enhancing model culture awareness (Li et al., 2024).
@@ 3 Collection of Culture-Gen
Table 1: Prompt for each Topic. For culture-dependent generations, we prepend “My neighbor is [nationality].” to the prompt.
@@
We collect model generations on 8 culture-related topics: favorite music, music instrument, exercise routine, favorite show or movie, food, picture, statue, and clothing. The task is to continue generating from natural language prompts shown in Table 1. To increase the compliance to topic, we add “Describe the [topic] of your neighbor.” in front of each prompt. We generate 100 samples for each culture on each topic.
@@ 4 Finding Culture Symbols in Culture-Gen
For llama2-13b and mistral-7b, we measure the association of a candidate symbol to a culture as the joint probability of Culture $c$ and Symbol $e$ conditioned on Topic $T$, written as $P(n,e|T)$  [^4] For mistral-7b, we perform a calibration process to mitigate the model’s prior bias towards a fixed set of cultures. See Appendix B.1.. More specifically, we construct a sentence following Table 5, and calculate the sentence probability using the same model that generated the candidate symbol. Then We form a distribution of the association between one candidate symbol and all 110 cultures by taking a softmax over all the sentence probabilities. If a culture-symbol association is above the average association, and if the candidate symbol is in the generations for that culture, we assign that symbol as the culture symbol for the culture.
@@ 5.2 Diversity of Culture Symbols: a measurement of LLM knowledge of cultural entities
Figure 5: Overlap in “music instrument”. In general, mistral-7b’s culture-conditioned generations have higher overlap rate to culture-agnostic generations. For both gpt-4 and mistral-7b, West European, English Speaking and Nordic countreis have the highest overlap rate.
@@ 6 Discussion
Pros and Cons of methodologies for assigning of culture symbols.
Naous et al. (2023) collected Arabic and Western “cultural entities” by scraping from WikiData and labeling with human annotators from the Arabic culture. While this approach obtains more accurate categorization of culture symbols, it is limited to the cultures from which high-quality human annotations are available, and thereby also limiting culture-bias research to such cultures. Our approach of automatically extracting culture symbol from cultural generations can encompass any culture of interest, although it currently cannot handle the boundary between “what is a possible answer for a culture” versus “what the people from that culture view as representative to their culture”. For example, in Culture-Gen “jogging” and “modern exercise” are generated and categorized as culture symbols for many cultures, as well as in culture-agnostic generations, even though such exercises may not be specific to any subset of cultures.
Simply asking a model to listing cultural symbols does not truly evaluate LM’s cultural perception.
@@ Appendix B Culture Symbols
B.1 Calibration of mistral-7b during unsupervised probability ranking.

### v4

@@ 1 Introduction
In this paper, we first present a simple LLM culture perception extraction framework that does not involve human labeling or external knowledge bases, and can be applied on any LLM and study any culture. We first use natural language prompts to elicit LLM generations on 8 culture-related topics for 110 countries and regions, using gpt-4, llama2-13b and mistral-7b. From these generation we extract culture symbols, i.e. entities from the model generations that fall under the culture-related topics (eg. “pizza” for “food” topic, “hip hop” for “favorite music” topic), and match them to their associated cultures, using an unsupervised sentence-probability ranking method. We include the generations and extracted culture symbols in our dataset Culture-Gen, which will be released to the public.
We then demonstrate how the community can use Culture-Gen for uncovering LLM’s global culture perceptions. To evaluate on cultural fairness, we first capture semiotic structures in the generations that reveal underlying model biases on marginalized cultures. We find that models tend to precede generations with the word “traditional” for cultures in Asian, Eastern European, and African-Islamic countries, as many as 30% for llama2-13b generations and almost 100% for gpt-4 generations. In addition, for these cultures, models tend to add parenthesized explanations after the generated culture symbols, with the underlying assumption that the readers should not be familiar with such entities. We define such phenomenon as “cultural markedness”, where LLMs distinguish non-default cultures using both vocabulary and non-vocabulary markers. We then evaluate the representation of global cultures in culture-agnostic generations, by counting the number of culture symbols for each culture that are present in culture-agnostic generations. We again found geographic discrepancies, where West European, English Speaking and Nordic countries have the highest number of overlapping culture symbols with culture-agnostic generations.
@@ 2 Related Work
Another line of work explores cultural diversity under multilingual settings (Mukherjee et al., 2023), incorporating geo-diverse multilingual probing (Yin et al., 2022), investigating multicultural biases using the Word Embedding Association Test (WEAT) across 24 languages (Mukherjee et al., 2023), and analyzing cross-lingual semantic alignments of cultural symbols through cultural proxies(Adilazuarda et al., 2024). In particular, low-resource languages have been studied as media of under-represented cultures like African American (Deas et al., 2023) and Indonesian Wibowo et al. (2023) languages and cultures. AlKhamissi et al. (2024). Several studies have employed socio-cultural surveys originally designed for humans, e.g., the World Values Survey (WVS) and Pew Global Attitudes Survey, to evaluate the cultural understanding of LLMs (Ramezani & Xu, 2023; Tao et al., 2023; Durmus et al., 2023) and extrapolate training data for enhancing model culture awareness (Li et al., 2024).
@@ 3 Collection of Culture-Gen
Table 1: Prompt for each Topic. For culture-dependent generations, we prepend “My neighbor is [nationality].” to the prompt. To increase the compliance to topic, we add “Describe the [topic] of your neighbor.” in front of each prompt.
@@
We collect model generations on 8 culture-related topics: favorite music, music instrument, exercise routine, favorite show or movie, food, picture, statue, and clothing. For topics of pictures and statues on the front door, we intended to extract culture symbols that do not belong to the other topics but still have cultural significance. For example, representative animals, religious symbols and historical figures of a culture can be extracted in this way. In downstream tasks settings such as story generation, one may ask the model to describe the picture, statue, decoration, etc. where the object depicted in these have cultural significance, but those objects may belong to different categories for different cultures. Therefore, we used “picture” and “statue” as an all-encompassing category, but what we were really interested in extracting are the objects that the model perceives to be depicted in those pictures and statues.
The task is to continue generating from natural language prompts shown in Table 1. We generate 100 samples for each culture on each topic.
@@ 4 Finding Culture Symbols in Culture-Gen
For llama2-13b and mistral-7b, we measure the association of a candidate symbol to a culture as the joint probability of Culture $c$ and Symbol $e$ conditioned on Topic $T$, written as $P(c,e|T)$  [^4] For mistral-7b, we perform a calibration process to mitigate the model’s prior bias towards a fixed set of cultures. See Appendix B.2.. More specifically, we construct a sentence following Table 5, and calculate the sentence probability using the same model that generated the candidate symbol. Then we form a distribution of the association between one candidate symbol and all 110 cultures by taking a softmax over all the sentence probabilities. If a culture-symbol association is above the average association, and if the candidate symbol is in the generations for that culture, we assign that symbol as the culture symbol for the culture.
@@
Ablation on the representativeness of Culture Symbols across demographics - age and gender.
We ablate on age using 17 years old and 70 years old, and on gender using male and female pronouns. The prompt and result is included in Appendix B.1.
@@ 5 LLM Global Culture Perception Analysis
To highlight the significance of markedness in culture-conditioned generations, we estimate the default presence of linguistic markers in Appendix C.
@@ 5.2 Diversity of Culture Symbols: a measurement of LLM knowledge of cultural entities
Figure 5: Overlap in “music instrument”. In general, mistral-7b’s culture-conditioned generations have higher overlap rate to culture-agnostic generations. For both gpt-4 and mistral-7b, West European, English Speaking and Nordic countries have the highest overlap rate.
@@ 6 Discussion
Our choice of continue-generation task for culture perception evaluation.
@@
For works that aim to collect comprehensive cultural data from LLMs and use it in downstream training or tuning language models, LLM-generated data may contain bias or hallucination. For more accurate categorization of culture symbols, works typically resort to online databases. For example,  Naous et al. (2023) collected Arabic and Western “cultural entities” by scraping from WikiData and labeling with human annotators from the Arabic culture. However, this approach is limited to the cultures from which high-quality human annotations are available, and thereby also limiting culture-bias research to such cultures. Since our work focuses on analyzing the cultural perceptions of the LLMs and does not intend the cultural symbols to be used in downstream training, the generations are allowed and intended to exhibit bias. Our approach of automatically extracting culture symbol from cultural generations can encompass any culture of interest, making it applicable in low-resource scenarios.
@@
Potential mitigation of uneven global cultural perception.
Our analyses of markedness, diversity of culture symbols, and default culture symbols show that current language models have uneven cultural perception and inadequate cultural knowledge, especially regarding marginalized cultures. Recent work suggests that such biases may have resulted from monotonic alignment procedures (Ryan et al., 2024). Therefore, we believe that mitigation of cultural biases lie in two steps: 1) Expanding the coverage of pretraining and instruction-tuning data to include global cultures so as to minimize undesirable behaviors in out-of-distribution settings, and 2) Pluralistic alignment (Sorensen et al., 2024), where models can recognize and generate according to diverse values and perspectives.
Sorensen et al. (2024) has suggested multiple potential directions for achieving pluralistic alignment, such as making models that output the whole spectrum of reasonable responses, models that can be faithfully steered to reflect certain properties or perspectives, and models whose distribution over answers matches that of a given target population.
In order to resolve conflicts between different cultures in culture-related generations, we believe that we can choose either of the directions above, depending on the intended usage of the model:
For general-purpose models, one may prefer models that output the whole spectrum of reasonable responses; for models developed with an audience in mind, such as personalized models, one may prefer models that can be steered for attributions preferred by that group; for data generation models, one may prefer models who can generate an array of answers during sampling.
@@ Appendix B Culture Symbols
B.1 Ablation Studies on Demographics
Hit rate
17 70 male female
exercise 48.3% (0.013) 41.3% (0.011) 55.9% (0.005) 52.4% (0.018)
food 51.9% (0.033) 56.9% (0.035) 62.0% (0.018) 63.6% (0.028)
music 42.4% (0.012) 37.9% (0.016) 45.0% (0.017) 42.2% (0.009)
New rate
17 70 male female
exercise 3.4% (0.0003) 5.7% (0.001) 6.8% (0.001) 6.4% (0.002)
food 5.5% (0.0004) 5.1% (0.0008) 6.3% (0.001) 6.0% (0.001)
music 22.8% (0.004) 16.3% (0.004) 21.0% (0.005) 18.5% (0.005)
Table 7: mistral-7b
Hit rate
17 70 male female
exercise 51.9% (0.038) 44.8% (0.029) 60.5% (0.034) 68.2% (0.168)
food 53.6% (0.015) 55.2% (0.023) 65.4% (0.015) 69.6% (0.021)
music 67.5% (0.014) 64.3% (0.016) 66.4% (0.003) 56.5% (0.010)
New rate
17 70 male female
exercise 4.3% (0.0005) 6.1% (0.0009) 9.5% (0.002) 6.4% (0.001)
food 3.4% (0.0002) 2.7% (0.0001) 5.5% (0.0003) 5.0% (0.0002)
music 18.6% (0.005) 8.2% (0.001) 13.7% (0.003) 11.7% (0.004)
Table 8: llama2-13b
For each model (llama2-13b and mistral-7b) and three representative topics (exercise routine, food, favorite music), we selected the culture with the most culture symbols extracted from each geographic region, totalling 12 cultures.
For studying age variance, we tried 17 and 70 years old, representing younger and older generations, and prompted with “Describe the [topic] of your neighbor. My neighbor is [culture] and is 17,70 years old. My neighbor probably …”. For studying gender variances, we prompted with “Describe the [topic] of your neighbor. My neighbor is [culture]. He, she probably …”
In Table 7 and 8, we report the mean and average of hit rate, percentage of culture symbols extracted from neutral prompts that are generated in each condition, and the mean and average of new rate, percentage of generated symbols that are not a culture symbol previously detected.
We find that in both gender and age conditions for both models, models only generate a part of the culture symbols extracted from the neutral generations (as reported by hit rate). This suggests that by increasing the number of conditions, the generated culture symbols become more homogenous. The rest of the generations are either culture symbols that belong to other cultures, or symbols not previously recognized as culture symbols (as reported by new rate). The new rate shows that for exercise routines and food, there’s only a small proportion of generations that have symbols unseen during neutral generations, and a slightly larger portion for music (understandable, as art genres differ more across subpopulations). We also would like to note that some of the newly discovered are not culture symbols, such as “another form of gentle exercise” in exercise routines of 70 year olds, but rather unique expressions, making the potentially unidentified culture symbols even fewer.
B.2 Calibration of mistral-7b during unsupervised probability ranking.
@@ Appendix C Markedness
Table 10 shows the number of vocabulary and parentheses markers in culture-agnostic generations for each topic and each model. Compared to markedness of culture-conditioned generations, the default degree of markedness on culture-related topics is very low.
favorite music exercise favorite show food picture statue clothing
music instrument routine or movie
llama2-13b vocab 0 2 1 0 0 0 0 0
paren 0 2 0 1 2 2 1 1
mistral-7b vocab 0 0 1 0 2 0 0 0
paren 2 0 1 3 3 2 2 3
gpt-4 vocab 0 0 0 0 10 1 1 0
paren 0 0 0 0 0 0 0 0
Table 10: Markedness in culture-agnostic generations for all models.